square root of (x+4)+4=square root of x+20
only x+4 and x+20 is square root

Found 2 solutions by mananth, jsmallt9:
Answer by mananth(16946)

sqrt(x+4)+4=square root of x+20
square both terms
..
(x+4)+4(sqrt(x+4)+16=x+20
x+20 +4(sqrt(x+4)=x+20
(-x -20)
4(sqrt(x+4))=0
/4
sqrt(x+4)=0
square
x+4 = 0
-4
x=-4
...
m.ananth@hotmail.ca

Answer by jsmallt9(3758)

I just solved this same problem for someone else. Click here to see the solution.